# Transformer-based stereo-aware 3D object detection from binocular images

### v3

@@ IV-C Comparison with existing methods
The performance of our TS3D on the KITTI test set is given in this subsection. The evaluation indicators of three difficulty levels (Easy, Moderate, and Hard) including 3D Car and Pedestrian detection tasks are displayed in the last row of Tab. V.
@@
As the first public Transformer-based binocular 3D object detector in the literature at the time of writing, TS3D significantly diminishes the performance gap between CNN-based and Transformer-based detectors. In the KITTI test set, the 3D Car and Pedestrian detection AP of TS3D respectively reaches 41.29% and 19.56%, which is competitive with that of the latest existing methods without LiDAR-based stereo matching supervision, i.e., YOLOStereo3D [5]. Compared with other models in the table, TS3D is based on the Transformer model. Compared with the first part of the table, where LiDAR-based stereo matching supervisions are used during training, TS3D does not rely on LiDAR to provide disparity annotations during training. Compared with other detectors that do not rely on LiDAR in the second part, in addition to the difference of the neural network model, the method of this paper has the following differences. From the perspective of multi-scale features, TLNet [42] and RT3DStereo [29] use single-scale feature maps; Whereas TS3D proposes SPFPN to generate multi-scale feature maps. Stereo R-CNN [30] only extracts multi-scale unary features, and uses single-scale features in the 3D object RoI head; TS3D uses SPFPN modules in stereo features, and extracts unary and stereo two pyramids of multi-scale features. RTS3D [31], YOLOStereo3D, and Disp R-CNN [26] also use the feature pyramid for both unary features and stereo features; but do not use a method similar to the SPFPN to fuse multi-scale stereo features. From the perspective of additional information required for training, in addition to 3D object annotations, TS3D requires a stereo matching algorithm to generate pseudo-ground-truth disparity maps; RT3DStereo relies on semantic segmentation annotations; RTS3D takes coarse 3D bounding boxes as input and gradually optimizes the detection results to achieve high-quality 3D object detection; Disp R-CNN needs to use external 3D models to generate more accurate pseudo point clouds for candidate objects. From the perspective of detection pipelines, Stereo R-CNN requires complex post-processing steps to obtain a more accurate 3D object bounding box after predicting the 3D bounding box; TS3D only uses 2D NMS [59] as the post-processing step. Stereo R-CNN, TLNet, and Disp R-CNN are two-stage algorithms [59]; whereas TS3D is a single-stage detector [60].
@@
Tab. V also lists the external dependencies and runtime of the existing methods that do not rely on LiDAR-based stereo matching supervisions. Among them, Disp R-CNN is higher than TS3D in Moderate 3D Car detection AP, however, its detection speed is slower, respectively 0.42 seconds and 0.67 seconds. TS3D takes only 0.09 seconds to detect from a pair of binocular images. In addition, TS3D merely relies on the pseudo-ground-truth of stereo matching, while Disp R-CNN requires instance segmentation supervision and 3D models during training; The peak GPU memory footprint of our TS3D during inference is 3.1 GB, which is memory efficient compared with 6.0 GB required by DSGN and 4.9 GB by LIGA-Stereo.
@@
We also visualize three detection results of TS3D on the KITTI validation set (Fig. 5, one column each). The first row visualizes the inverse-projection of disparity map predicted by TS3D and detected 3D bounding boxes (pink); The second row illustrates the projection of detected 3D bounding boxes onto the left image. The last row shows the projection of 3D ground-truth boxes on the left image. In the first column in Fig. 5, the shadow of trees on the side of the road is a challenge for image-based object detectors, and the distant vehicle ahead is of relatively small-scale. Thanks to the explicitly embedded rich details and correspondence information in multi-scale stereo features and Transformer decoders, TS3D can alleviate the problems of scale variation and shadow defects. The second column shows a roadside parking scene on a narrow road. In addition to the detected roadside cars of various scales, TS3D is also able to detect the car parked vertically on the left. The third column showcases the detection of distant cars. All four cars are detected, however, the right two cars are of erroneous orientations due to their small scale.
@@ V Conclusions
Transformer is making progress in various fields of computer vision and is expected to be compatible with binocular 3D object detection. We have proposed a Transformer-based Stereo-aware 3D object detector (TS3D), which has successfully applied the Transformer model to the binocular 3D object detection task by designing Disparity-Aware Positional Encoding (DAPE) and Stereo Preserving Feature Pyramid Network (SPFPN). The image correspondence and 2D location information have been explicitly encoded into object queries by DAPE, providing the decoder with the spatial information of the 3D scene. The SPFPN has been designed to provide the Transformer decoder with enriched multi-scale features and object embeddings. In the fusion process, SPFPN has preserved the image correspondence information in the stereo features; In the cross-scale aggregation, the complementary image correspondence information across scales has been aggregated, and eventually, the stereo feature pyramid with enriched details and correspondence information has been obtained.
@@
Acknowledgments
This work was partially supported bythe National Key R&D Program of China under Grant 2022ZD0160400, the National Key R&D Program of China under Grant 2018AAA0102800, the Tianjin Science and Technology Program under Grant 19ZXZNGX00050, and the Tianjin Natural Science Foundation under Grant JCQNJC00420.

### v4

@@ IV-B3 The number of decoder layers and the usage of intermediate supervision
IV-B4 Discussion on performance drops of existing surround-view Transformer detectors
Broadly speaking, a surround-view scene is of large field-of-view (FOV) and adjacent views are with small overlapping areas; A classic binocular stereo scene is of relatively smaller FOV and the overlapping areas are large, which enables image correspondence to represent more precise 3D information. (1) Adapting surround-view detectors to the binocular scene narrows the FOV and introduces large overlaps. However, the former results in less 3D scene information, and the latter is not fully considered in existing surround-view detectors. As a consequence, the detection precision significantly drops (shown in Fig. 1b). We tackle that problem by introducing image correspondence representations to the Transformer model. (2) Another reason for the performance drops is insufficient training samples. Existing Transformer-based multi-view 3D object detectors are trained with 1.4 million object annotations on the nuScenes dataset. However, the binocular dataset, KITTI, contains merely $1/7$ object annotations of those in nuScenes. In our TS3D, that problem is alleviated by the NPAQuery.
@@ IV-C Comparison with existing methods
The performance of our TS3D on the KITTI test set is given in this subsection. The evaluation indicators of three difficulty levels (Easy, Moderate, Hard) including 3D Car and Pedestrian detection tasks are displayed in the last row of Tab. V.
@@
As the first public Transformer-based binocular 3D object detector in the literature at the time of writing, TS3D significantly diminishes the performance gap between CNN-based and Transformer-based detectors. In the KITTI test set, the 3D Car and Pedestrian detection AP of TS3D respectively reaches 41.29% and 19.56%, which is competitive with that of the latest existing methods without LiDAR-based stereo matching supervision, i.e., YOLOStereo3D [5]. Compared with other models in the table, TS3D is based on the Transformer model. Compared with the first part of the table, where LiDAR-based stereo matching supervisions are used during training, TS3D does not rely on LiDAR to provide disparity annotations during training. Compared with other detectors that do not rely on LiDAR in the second part, in addition to the difference in the neural network model, the method of this paper has the following differences. From the perspective of multi-scale features, TLNet [42] and RT3DStereo [29] use single-scale feature maps; Whereas TS3D proposes SPFPN to generate multi-scale feature maps. Stereo R-CNN [30] only extracts multi-scale unary features, and uses single-scale features in the 3D object RoI head; TS3D uses SPFPN modules in stereo features, and extracts unary and stereo two pyramids of multi-scale features. RTS3D [31], YOLOStereo3D, and Disp R-CNN [26] also use the feature pyramid for both unary features and stereo features; but do not use a method similar to the SPFPN to fuse multi-scale stereo features. From the perspective of additional information required for training, in addition to 3D object annotations, TS3D requires a stereo matching algorithm to generate pseudo-ground-truth disparity maps; RT3DStereo relies on semantic segmentation annotations; RTS3D takes coarse 3D bounding boxes as input and gradually optimizes the detection results to achieve high-quality 3D object detection; Disp R-CNN needs to use external 3D models to generate more accurate pseudo point clouds for candidate objects. From the perspective of detection pipelines, Stereo R-CNN requires complex post-processing steps to obtain a more accurate 3D object bounding box after predicting the 3D bounding box; TS3D only uses 2D NMS [59] as the post-processing step. Stereo R-CNN, TLNet, and Disp R-CNN are two-stage algorithms [59]; whereas TS3D is a single-stage detector [60].
@@
Tab. V also lists the external dependencies and runtime of the existing methods that do not rely on LiDAR-based stereo matching supervisions. Among them, Disp R-CNN is higher than TS3D in Moderate 3D Car detection AP, however, its detection speed is slower, respectively 0.42 seconds and 0.67 seconds. TS3D takes only 0.09 seconds to detect from a pair of binocular images. In addition, TS3D merely relies on the pseudo-ground-truth of stereo matching, while Disp R-CNN requires instance segmentation supervision and 3D models during training; The peak GPU memory footprint of our TS3D during inference is 3.1 GB, which is memory efficient compared with the 6.0 GB required by DSGN and 4.9 GB by LIGA-Stereo.
@@
We also visualize three detection results of TS3D on the KITTI validation set (Fig. 5, one column each). The first row visualizes the inverse projection of the disparity map predicted by TS3D and detected 3D bounding boxes (pink); The second row illustrates the projection of detected 3D bounding boxes onto the left image. The last row shows the projection of 3D ground-truth boxes on the left image. In the first column in Fig. 5, the shadow of trees on the side of the road is a challenge for image-based object detectors, and the distant vehicle ahead is of a relatively small scale. Thanks to the explicitly embedded rich details and correspondence information in multi-scale stereo features and Transformer decoders, TS3D can alleviate the problems of scale variation and shadow defects. The second column shows a roadside parking scene on a narrow road. In addition to the detected roadside cars of various scales, TS3D is also able to detect the car parked vertically on the left. The third column showcases the detection of distant cars. All four cars are detected, however, the right two cars are of erroneous orientations due to their small scale.
@@ V Conclusions
Transformers are making progress in various fields of computer vision and are expected to be compatible with binocular 3D object detection. We have proposed a Transformer-based Stereo-aware 3D object detector (TS3D), which has successfully applied the Transformer model to the binocular 3D object detection task by designing Disparity-Aware Positional Encoding (DAPE) and Stereo Preserving Feature Pyramid Network (SPFPN). The image correspondence and 2D location information have been explicitly encoded into object queries by DAPE, providing the decoder with the spatial information of the 3D scene. The SPFPN has been designed to provide the Transformer decoder with enriched multi-scale features and object embeddings. In the fusion process, SPFPN has preserved the image correspondence information in the stereo features; In the cross-scale aggregation, the complementary image correspondence information across scales has been aggregated, and eventually, the stereo feature pyramid with enriched details and correspondence information has been obtained.